OpenStudy (abbycross167):

Can someone please show me the steps to answer this question? The width, w, of a rectangular rug is 4 less than its length, L. If the area of the rug is 120 square feet, which equation could be used to find the dimensions of the rug?

OpenStudy (instagrammodel):

I'll give you an example of one, Hopefully this will help you a bit. :) w(2w+3) = 324 2w^2+3w-324 = 0 (2w+27)(w-12) = 0 w=12, so the room is 12x27

OpenStudy (abbycross167):

@InstagramModel I'm sorry I forgot to put my answer choices on here a. L(4-L) = 120 b. L(L-4) = 120 c. 2(L-4) + 2L = 120 d. 2w + 2L = 120

OpenStudy (abbycross167):

I can't really figure out what the steps are...

OpenStudy (anonymous):

d. 2w + 2L = 120

OpenStudy (anonymous):

i took the assesment

OpenStudy (abbycross167):

Did you really? what school do you go to?